OpenStudy (caozeyuan):

How can I convert the polar equation r=1-cos(theta) into its cartesian equivalent? How about r=1+sin(theta) and 2/r=4+cos(theta)?

OpenStudy (caozeyuan):

@amistre64 Help me!!!

OpenStudy (caozeyuan):

I'm dying! some one save me please!!!!

OpenStudy (amistre64):

some things to keep in mind x = r cos(t) y = r sin(t) r^2 = x^2+y^2

OpenStudy (caozeyuan):

I just got a result. For the first one, it's x^2+y^2=sqrt(x^2-y^2)-y. Is that correct

OpenStudy (amistre64):

the polar plot of r = 1-cos(t) is heart shaped x^2 + y^2 = r - rcos(t) x^2 + y^2 = r - x x^2 - x + y^2 = r (x-1/2)^2 + y^2 - 1/4 = r yeah, these things have always been a pain to me too

OpenStudy (amistre64):

but your solution so far is not quite correct yet http://www.wolframalpha.com/input/?i=x%5E2%2By%5E2%3Dsqrt%28x%5E2-y%5E2%29-y

OpenStudy (amistre64):

my getup is pointing in the wrong direction, but is getting closer :) http://www.wolframalpha.com/input/?i=%28x-1%2F2%29%5E2%2By%5E2-1%2F4%3Dsqrt%28x%5E2%2By%5E2%29

OpenStudy (amistre64):

if i change the (x-1/2) into (-x-1/2) we get a better result

OpenStudy (amistre64):

pfft, i see what i did to begin with, i should have add x to each side, that where i introduced my error at

OpenStudy (caozeyuan):

you see. r=1-cos theta, r^2 = r-rcos theta. x^2+y^2=r-y, x^2+y^2=sqrt(x^2+y^2)-y why, is my reasoning wrong?

OpenStudy (caozeyuan):

@amistre64

OpenStudy (amistre64):

because rcos(t) = x, not y

OpenStudy (amistre64):

otherwise, your spot on

OpenStudy (caozeyuan):

I'sooooo stupid!